# revtex4-2Repair the float

### v3

@@ Appendix B Numerical set up for the Diagonalization of the Bogoliubov-de Gennes Equations
This section outlines the set up for numerical diagonalization of the BdG Hamiltonian introduced in the last section. When we are considering a singlet SC, we can simplify the superconducting matrix to the following form
$\mathbf{\tilde{\Delta}}=\left[\begin{array}[]{c c c c}\mathbf{0}&\mathbf{\Delta}\\ -\mathbf{\Delta}&\mathbf{0}\end{array}\right],\;\;\mathbf{\Delta}^{T}=\mathbf{\Delta}$ (B.1)
@@
$\Psi=\begin{bmatrix}\mathbf{c}_{\uparrow}&\mathbf{c}^{\dagger}_{\downarrow}&\mathbf{c}_{\uparrow}&\mathbf{c}^{\dagger}_{\uparrow}&\end{bmatrix}^{T}=\mathbf{O}\tilde{\Psi}\Rightarrow$ $\hat{H}=\Psi^{\dagger}\mathbf{V}\mathbf{E}\mathbf{V}^{\dagger}\Psi$ (B.2)
The orthogonal transformation in question takes the following simple form and gives an explicit relation between the similarity transforms in question
$\mathbf{O}=\begin{bmatrix}\mathbf{I}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}\\ \mathbf{0}&\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{I}&\mathbf{0}\end{bmatrix},\;\mathbf{\tilde{V}}=\mathbf{O}^{T}\mathbf{V}$ (B.3)
Defining the following matrices
$\mathbf{H}_{\pm}=\begin{bmatrix}\mathbf{t}&\pm\mathbf{\Delta}\\ \pm\mathbf{\Delta}^{\dagger}&-\mathbf{t}\end{bmatrix}$ (B.4)
we find that our transformed Hamiltonian takes the following form ($\psi_{\pm}$ are defined inline according to equation (B.2))
@@
The matrix is block diagonal, meaning we can diagonalize the two sub-Hamiltonians independently. The eigenvectors of the lower block are related to those of the upper via the same transformation shown in appendix A equation (A.5). In addition to this, the energies of each block come in $\pm$ pairs. This can be seen by complex conjugating the eigenvalue equations defined by equation (B.4), then applying the following orthogonal matrix: $\mathbf{O}^{\prime}=\begin{bmatrix}-\mathbf{I}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}\end{bmatrix}$.
Solving one of these sub-blocks is enough. After doing so, and organizing our eigenvectors from least to greatest (in energy), we arrive at the following similarity transforms
$\mathbf{V}_{+}$ $=\begin{bmatrix}\mathbf{v}^{*}&\mathbf{u}\\ -\mathbf{u}^{*}&\mathbf{v}\end{bmatrix},\;\;\mathbf{V}_{-}=\begin{bmatrix}\mathbf{v}^{*}&-\mathbf{u}\\ \mathbf{u}^{*}&\mathbf{v}\end{bmatrix}\;\;\Rightarrow\;\;\mathbf{V}=\begin{bmatrix}\mathbf{V}_{+}&\mathbf{0}\\ \mathbf{0}&\mathbf{V}_{-}\end{bmatrix}$ (B.6)
Thus, we can simply diagonalize $\mathbf{H}_{+}$, then use our above string of relations to find the coherence factors given in the previous appendix if needed. We work with these coherence factors defined above in the following section.
@@ Appendix C Retarded Green Function at zero Temperature
Using the results laid out in the previous sections we can find the retarded zero temperature Green functions. We start with the real space retarded Green function at zero temperature
@@
$=$ $-i\theta(t)\big{\langle}\big{\{}e^{i\hat{H}t}\hat{c}_{\mathbf{r}_{i}\sigma}e^{-i\hat{H}t},\hat{c}^{\dagger}_{\mathbf{r}_{j}\sigma}\big{\}}\big{\rangle}$
Using the action of our quasi-particle operators on our ground state, and relation (equation (A.7) or equation (B.6)) we can evaluate these terms. Recall we are in the singlet configuration, so we may work with a single spin, say, $\sigma=\uparrow$, and drop the spin label:
$G(\mathbf{r}_{i},\mathbf{r}_{j},t)$ $=-i\theta(t)\sum_{E_{l}>0}\Big{(}v_{il}v_{jl}^{*}e^{iE_{l}t}+u^{*}_{il}u_{jl}e^{-iE_{l}t}\Big{)}$
We are interested in the Fourier transform which takes us from the time domain to the frequency domain. This integral requires a dampening factor, $\epsilon$, for convergence, which represents the energy resolution, taken to be 2.5 meV. The Fourier transform gives us the Green function in a familiar form
@@
We’ll also need to Fourier Transform to $k$-space to obtain the spectral function
$G(\mathbf{k_{a}},\mathbf{k_{b}},\omega)$ $=\frac{1}{N}\sum_{i,j}G(\mathbf{r}_{i},\mathbf{r}_{j},\omega)e^{i\mathbf{k_{a}}\cdot\mathbf{r_{i}}}e^{-i\mathbf{k_{b}}\cdot\mathbf{r_{j}}}=\sum_{E_{l}\geq 0}\Bigg{(}\frac{\tilde{u}_{l}(\mathbf{k_{a}})\tilde{u}^{*}_{l}(\mathbf{k_{b}})}{\omega-E_{l}+i\epsilon}+\frac{\tilde{v}_{l}(\mathbf{k_{b}})\tilde{v}^{*}_{l}(\mathbf{k_{a}})}{\omega+E_{l}+i\epsilon}\Bigg{)}$ (C.2)
@@
$\tilde{u}_{l}(\mathbf{k})$ $=\frac{1}{N}\sum_{i}u_{il}e^{i\mathbf{k}\cdot\mathbf{r_{i}}},\;\;\;\tilde{v}_{l}(\mathbf{k})=\frac{1}{N}\sum_{i}v_{il}e^{i\mathbf{k}\cdot\mathbf{r_{i}}}$
Figure 12: (Online Color) Comparison plots of the LDOS for a PDW order SC with the three topological defects (described in equation (4.2).) First, we consider the zero bias LDOS given in the first row for (a) the half-vortex, (b) the Abrikosov vortex and (c) the double dislocation. In the second row we include plots of the LDOS at probing voltages of .$75\bar{\Delta}.$ for (d) the half-vortex, (e) the Abrikosov vortex and (f) the double dislocation. Each plot is normalized to the scale of the defect as in the main text.
Figure 13: (Online Color) The Fourier transform of the LDOS for an Abrikosov vortex of the PDW state for a range of probing voltages. We take cuts along the line $k_{y}=0$ and plot the absolute value of the FT vs probing voltage. This figure showcases the QPI for the Abrikosov vortex, to which we note the dispersing quasi-particles; hence, these patterns do not represent a bound state. We suppressed the intensity around the $\Gamma$-point in this plot. A similar pattern of the QPI is present in the case of the double dislocation suggesting it too does not posses bound quasi-particles.
@@ Appendix D Plots of $\rho_{2\mathbf{Q}}$ and the LDOS
This appendix provides comparison plots between the induced CDW for our three defects using equation (2.6) directly. We still see the signatures of the defects described in the text (e.g. a edge dislocation). Representative plots of the $\rho_{2\bm{Q}}$ are shown in figure 3. Comparison plots to the LDOS at zero bias is shown in figures 12a-12c. Note the core of the vortex is noticeably different in figure 3 than in figure 12 for the subplots which correspond to the full-vortex and the double dislocation. This is due to the fact the LDOS calculation possesses information regarding the quasi-particles and the patterns are a result of quasi-particle-interference. The half-vortex looks similar in both these figures because its core contains a fully gapped FF state. On the other hand, the other two defects, the full-vortex and the double dislocation, possess order parameters with a vanishing superconducting gap at the core. The integrated LDOS (the charge density) for the full-vortex and the double dislocation also possesses a nonzero weight within the vortex core due to the fact the vortex can accommodated quasi-particles, but the dynamic features are integrated out. In figures 12d-12f we also provide a set of LDOS at a bias of .$75\bar{\Delta}.$. Notice the depletion of states that occurs in the core of the vortex even in the case of a half-vortex where there is a partial gap present.
@@
Note these quasi-particles do not reside solely in the vortex core and are thus not bound states. We can illustrate this by taking a FT of the LDOS and plotting cuts in momentum space over a range of energies to observe a dispersion relation. The nonzero Fourier harmonics at each energy, $\omega$, corresponds to the scattering wave vectors connecting different regions of the surface determined by the spectral function evaluated at the same $\omega$. The regions of large joint-DOS on this surface provides us with the dominate Fourier harmonics, and if these regions disperse, we see it in the FT of the LDOS. In figure 13 the QPI of the Abrikosov vortex is provided for cuts in $k$-space along the nodal direction, and it indicates we have dispersing quasi-particles by the change in Fourier harmonics.
@@
Figure 14: Comparison of $\frac{dI}{dV}$ curves related to the tunneling DOS for the half-vortex (taken to live on the CuO-bonds and not the center of the plaquette) at different lattice sites, indicated with the $x$ position relative to the vortex core (with a relative shift of 1 between curves). We look at two sets of curves taken outside the vortex core of radius $r_{0}=24a_{0}$, (a) one to the left (negative $x$) and (b) one to the right (positive $x$). Both sets of curves repeat every four lattice spacings (in $x$), like an LO state. This indicates a relative shift between the curves in the two columns by two lattice spacings. The shift seen in these plots are a result of the jump in $\theta_{+}$, defined in equation (2.8), occurring across the core of the vortex.
@@
Figure 15: Comparison of $\frac{dI}{dV}$ curves related to the tunneling DOS for the half-vortex, the double dislocation and the Abrikosov vortex at two different lattice sites. The amplitudes are expressed in arbitrary units, each being normalized to the scale of the respective defect. In terms of fractions of the halo radius, $r_{0}=24a_{0}$, the plots are taken at a distance of (a) .$025\cdot r_{0}.$ and (b) $4\cdot r_{0}$ in respect to the vortex cores. Inside the core (a), the double dislocation and the Abrikosov vortex resemble a free particle dispersion since both the PDW components are small here. However, in the case of the half-vortex the core state is still superconducting, resembling a squeezed in FF state. Outside the core (b) all the curves resemble the pure PDW, but the half-vortex possesses additional satellite peaks due to the fact it possesses inequivalent PDW components.
@@ Appendix E Tunneling DOS Comparison
Here we compare the plots corresponding to the tunneling DOS for the half-vortex outside the vortex core on its LHS and RHS (figures 14a and 14b respectively). The tunneling DOS are labeled with only $x$-coordinates to indicate how far we are in respect to the half-vortex core ($x$ being measured relative to its center). This indicates we are looking at consecutive $x$-positions in a given sub figures as well as we are comparing curves to the left (negative sign) and right (positive sign) of the vortex. The particular positions used for the representative curves are somewhat arbitrary in the sense we only want positions outside the core. Also, note the defect is placed on the CuO-bonds here opposed to the center of the plaquette as was done in the text for the LDOS. The spectral functions corresponding to these two situations are equivalent, meaning the energetics are very similar, but the shape of the $dI/dV$ curves will certainly change character since the gap amplitude takes on different values on the lattice sites.
@@
Both set of curves, to the right and to the left of the vortex core, have a periodicity of $4$ lattice spacings, just like that of the pure PDW in the text (the same conclusion holds for the placement of the defect at the center of the plaquette). However, unlike a pure LO state the data appears to have a relative shift of $2$ lattice spacings when comparing the curves from the right to the left (again this holds true for the half-vortex placed at the center of the plaquette). As was discussed in the text, this is due to the jump in $\theta_{+}$ by $\pi/2$ across the half-vortex’s core and can be explained as an accumulation in phase due to the non-zero COM-momentum of our Cooper pairs. That is: $\mathbf{Q}\cdot\mathbf{r}=\pi/2$ if $\mathbf{r}=2a_{0}\mathbf{e}_{x}$. This is a smoking gun signature of the half-vortex since a bona-fied LO state would not break inversion symmetry like so.
@@
Now we compare the tunneling DOS of our topological defects in and outside the core of the vortex (figures 15a and 15b, respectively). The Abrikosov vortex and the double dislocation both resemble a free particle tunneling DOS, albeit some additional wiggles. This is simply because we are not right at the center of the core. Also, notice the curves for these two defects are not in perfect agreement here, due to the form of the topological defect. Indeed, the electron DOS has to be distinct in the core of each vortex due to the difference in the respective defect. The half-vortex still possesses signatures of superconductivity in the core due to the fact it has FF character there. In fact, it resembles a squeezed in FF-state.
@@
Outside the core the double dislocation and the Abrikosov vortex look almost identical to one another and also the pure PDW (LO-state). This is because the topological defect has less of an influence on the electronic states far away from the core. The half-vortex has additional wiggles in its tunneling DOS corresponding to satellite peaks. These additional peaks arise here and not for the other two defects because of the half-vortex’s inequivalent Fourier components.
@@
Figure 16: Comparison plots of various Fermi surfaces. First, we provide a plot for (a) a defect-free PDW with an s-wave form factor compared with (b) a defect-free PDW with a d-wave form factor superimposed with the normal state dispersion (shown in light grey). The main arcs seen in both plots correspond to the normal state dispersion. In (c) we have a log plot of the Abrikosov vortex with a d-wave form factor, which greatly resembles the double dislocation and the pure PDW. Both defects prompt a slight redistribution of spectral weight, but nothing too dramatic like the half-vortex. Weaker features are uncovered with the log plots where the additional arcs correspond to the normal state dispersion shifted by ±$\pm\mathbf{Q}$.
@@ Appendix F Comparison of Spectral functions for s-wave and d-wave form factors and additional order parameters
In this section we compare some additional spectral functions not included in the main text. First, let’s focus in on the spectral functions corresponding to defect-free PDW states, one with an s-wave form factor the other d-wave (figures 16a and 16b, respectively). The normal state dispersion is retained in both cases along the so called Fermi arcs, so certain features are robust. The LDOS looks more or less the same for both form factors (not shown). Figures 16c is the natural log of the spectral function corresponding to the Abrikosov vortex, which closely resembles the double dislocation (not shown). Slight re-distributions of the spectral weights, occur for both these defects, but nothing as dramatic as that seen in the case of the half-vortex.

### v4

@@ Appendix B Numerical set up for the Diagonalization of the Bogoliubov-de Gennes Equations
This appendix outlines the set up for numerical diagonalization of the BdG Hamiltonian introduced in the last appendix. When we are considering a singlet SC, we can simplify the superconducting matrix to the following form:
$\mathbf{\tilde{\Delta}}=\left[\begin{array}[]{c c c c}\mathbf{0}&\mathbf{\Delta}\\ -\mathbf{\Delta}&\mathbf{0}\end{array}\right],\;\;\mathbf{\Delta}^{T}=\mathbf{\Delta}.$ (B.1)
@@
$\Psi=\begin{bmatrix}\mathbf{c}_{\uparrow}&\mathbf{c}^{\dagger}_{\downarrow}&\mathbf{c}_{\uparrow}&\mathbf{c}^{\dagger}_{\uparrow}&\end{bmatrix}^{T}=\mathbf{O}\tilde{\Psi}\Rightarrow$ $\hat{H}=\Psi^{\dagger}\mathbf{V}\mathbf{E}\mathbf{V}^{\dagger}\Psi.$ (B.2)
The orthogonal transformation in question takes the following simple form and gives an explicit relation between the similarity transforms in question:
$\mathbf{O}=\begin{bmatrix}\mathbf{I}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}\\ \mathbf{0}&\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{I}&\mathbf{0}\end{bmatrix},\;\mathbf{\tilde{V}}=\mathbf{O}^{T}\mathbf{V}.$ (B.3)
Defining the following matrices:
$\mathbf{H}_{\pm}=\begin{bmatrix}\mathbf{t}&\pm\mathbf{\Delta}\\ \pm\mathbf{\Delta}^{\dagger}&-\mathbf{t}\end{bmatrix},$ (B.4)
we find that our transformed Hamiltonian takes the following form [$\psi_{\pm}$ are defined inline according to Eq.  (B.2)]:
@@
The matrix is block diagonal, meaning we can diagonalize the two sub-Hamiltonians independently. The eigenvectors of the lower block are related to those of the upper via the same transformation shown in Eq.  (A.5). In addition to this, the energies of each block come in $\pm$ pairs. This can be seen by complex conjugating the eigenvalue equations defined by Eq.  (B.4), then applying the following orthogonal matrix: $\mathbf{O}^{\prime}=\begin{bmatrix}\mathbf{0}&\mathbf{I}\\ -\mathbf{I}&\mathbf{0}\end{bmatrix}$.
Solving one of these sub-blocks is enough. After doing so, and organizing our eigenvectors from least to greatest (in energy), we arrive at the following similarity transforms:
$\mathbf{V}_{+}$ $=\begin{bmatrix}\mathbf{v}^{*}&\mathbf{u}\\ -\mathbf{u}^{*}&\mathbf{v}\end{bmatrix},\;\;\mathbf{V}_{-}=\begin{bmatrix}\mathbf{v}^{*}&-\mathbf{u}\\ \mathbf{u}^{*}&\mathbf{v}\end{bmatrix}\;\;\Rightarrow\;\;\mathbf{V}=\begin{bmatrix}\mathbf{V}_{+}&\mathbf{0}\\ \mathbf{0}&\mathbf{V}_{-}\end{bmatrix}.$ (B.6)
Thus, we can simply diagonalize $\mathbf{H}_{+}$, then use our above string of relations to find the coherence factors given in the previous appendix if needed. We work with these coherence factors defined above in the following appendix.
@@ Appendix C Retarded Green Function at zero Temperature
Using the results laid out in the previous appendixes we can find the retarded zero temperature Green functions. We start with the real space retarded Green function at zero temperature
@@
$=$ $-i\theta(t)\big{\langle}\big{\{}e^{i\hat{H}t}\hat{c}_{\mathbf{r}_{i}\sigma}e^{-i\hat{H}t},\hat{c}^{\dagger}_{\mathbf{r}_{j}\sigma}\big{\}}\big{\rangle}.$
Using the action of our quasiparticle operators on our ground state, and relation (Eq.  (A.7) or Eq.  (B.6)), we can evaluate these terms. Recall that we are in the singlet configuration, so we may work with a single spin, say, $\sigma=\uparrow$, and drop the spin label:
$G(\mathbf{r}_{i},\mathbf{r}_{j},t)$ $=-i\theta(t)\sum_{E_{l}>0}\Big{(}v_{il}v_{jl}^{*}e^{iE_{l}t}+u^{*}_{il}u_{jl}e^{-iE_{l}t}\Big{)}.$
We are interested in the Fourier transform, which takes us from the time domain to the frequency domain. This integral requires a dampening factor, $\epsilon$, for convergence, which represents the energy resolution, taken to be 2.5 meV. The Fourier transform gives us the Green function in a familiar form:
@@
We’ll also need to Fourier transform, to $k$-space to obtain the spectral function
$G(\mathbf{k_{a}},\mathbf{k_{b}},\omega)$ $=\frac{1}{N}\sum_{i,j}G(\mathbf{r}_{i},\mathbf{r}_{j},\omega)e^{i\mathbf{k_{a}}\cdot\mathbf{r_{i}}}e^{-i\mathbf{k_{b}}\cdot\mathbf{r_{j}}}=\sum_{E_{l}\geq 0}\Bigg{(}\frac{\tilde{u}_{l}(\mathbf{k_{a}})\tilde{u}^{*}_{l}(\mathbf{k_{b}})}{\omega-E_{l}+i\epsilon}+\frac{\tilde{v}_{l}(\mathbf{k_{b}})\tilde{v}^{*}_{l}(\mathbf{k_{a}})}{\omega+E_{l}+i\epsilon}\Bigg{)}.$ (C.2)
@@
$\tilde{u}_{l}(\mathbf{k})$ $=\frac{1}{N}\sum_{i}u_{il}e^{i\mathbf{k}\cdot\mathbf{r_{i}}},\;\;\;\tilde{v}_{l}(\mathbf{k})=\frac{1}{N}\sum_{i}v_{il}e^{i\mathbf{k}\cdot\mathbf{r_{i}}}.$
Figure 12: (Online Color) Comparison plots of the LDOS for a PDW order SC with the three topological defects [described in Eq.  (4.2)]. First, we consider the zero bias LDOS given in the first row for (a) the half-vortex, (b) the Abrikosov vortex and (c) the double dislocation. In the second row we include plots of the LDOS at probing voltages of $0.75\bar{\Delta}$ for (d) the half-vortex, (e) the Abrikosov vortex, and (f) the double dislocation. Each plot is normalized to the scale of the defect as in the main text.
Figure 13: (Online Color) The Fourier transform of the LDOS for an Abrikosov vortex of the PDW state for a range of probing voltages. We take cuts along the line $k_{y}=0$ and plot the absolute value of the FT vs probing voltage. This figure showcases the QPI for the Abrikosov vortex, for which we note the dispersing quasiparticles; hence, these patterns do not represent a bound state. We suppressed the intensity around the $\Gamma$-point in this plot. A similar pattern of the QPI is present in the case of the double dislocation, suggesting it too does not possess bound quasiparticles.
@@ Appendix D Plots of $\rho_{2\mathbf{Q}}$ and the LDOS
This appendix provides comparison plots between the induced CDW for our three defects using Eq.  (2.6) directly. We still see the signatures of the defects described in the text (e.g., a edge dislocation). Representative plots of the $\rho_{2\bm{Q}}$ are shown in Fig. 3. Comparison plots to the LDOS at zero bias are shown in Figs. 12a-12c. Note that the core of the vortex is noticeably different in Fig. 3 than in Fig. 12 for the subplots which correspond to the full vortex and the double dislocation. This is due to the fact the LDOS calculation possesses information regarding the quasiparticles, and the patterns are a result of quasiparticle interference. The half-vortex looks similar in both these figures because its core contains a fully gapped FF state. On the other hand, the other two defects, the full vortex and the double dislocation, possess order parameters with a vanishing superconducting gap at the core. The integrated LDOS (the charge density) for the full vortex and the double dislocation also possesses a nonzero weight within the vortex core due to the fact the vortex can accommodated quasiparticles, but the dynamic features are integrated out. In Figs. 12d-12f we also provide a set of LDOS at a bias of $0.75\bar{\Delta}$. Notice the depletion of states that occurs in the core of the vortex even in the case of a half-vortex where there is a partial gap present.
@@
Note these quasiparticles do not reside solely in the vortex core and are thus not bound states. We can illustrate this by taking a FT of the LDOS and plotting cuts in momentum space over a range of energies to observe a dispersion relation. The nonzero Fourier harmonics at each energy, $\omega$, corresponds to the scattering wave vectors connecting different regions of the surface determined by the spectral function evaluated at the same $\omega$. The regions of large joint DOS on this surface provides us with the dominate Fourier harmonics, and if these regions disperse, we see it in the FT of the LDOS. In Fig. 13 the QPI of the Abrikosov vortex is provided for cuts in $k$-space along the nodal direction, and it indicates we have dispersing quasiparticles by the change in Fourier harmonics.
@@
Figure 14: Comparison of $\frac{dI}{dV}$ curves related to the tunneling DOS for the half-vortex (taken to live on the CuO bonds and not the center of the plaquette) at different lattice sites, indicated with the $x$ position relative to the vortex core (with a relative shift of 1 between curves). We look at two sets of curves taken outside the vortex core of radius $r_{0}=24a_{0}$, (a) one to the left (negative $x$) and (b) one to the right (positive $x$). Both sets of curves repeat every four lattice spacings (in $x$), like an LO state. This indicates a relative shift between the curves in the two columns by two lattice spacings. The shift seen in these plots are a result of the jump in $\theta_{+}$, defined in Eq.  (2.8), occurring across the core of the vortex.
@@
Figure 15: Comparison of $\frac{dI}{dV}$ curves related to the tunneling DOS for the half-vortex, the double dislocation and the Abrikosov vortex at two different lattice sites. The amplitudes are expressed in arbitrary units, each being normalized to the scale of the respective defect. In terms of fractions of the halo radius, $r_{0}=24a_{0}$, the plots are taken at a distance of (a) $0.025\;r_{0}$ and (b) $4\;r_{0}$ in respect to the vortex cores. Inside the core (a), the double dislocation and the Abrikosov vortex resemble a free particle dispersion since both the PDW components are small here. However, in the case of the half-vortex the core state is still superconducting, resembling a squeezed in FF state. Outside the core (b) all the curves resemble the pure PDW, but the half-vortex possesses additional satellite peaks due to the fact it possesses inequivalent PDW components.
@@ Appendix E Tunneling DOS Comparison
Here we compare the plots corresponding to the tunneling DOS for the half-vortex outside the vortex core on its l.h.s. and r.h.s. [Figs. 14a and 14b respectively]. The tunneling DOS are labeled with only $x$ coordinates to indicate how far we are in respect to the half-vortex core ($x$ being measured relative to its center). This indicates we are looking at consecutive $x$ positions in a given sub figures as well as we are comparing curves to the left (negative sign) and right (positive sign) of the vortex. The particular positions used for the representative curves are somewhat arbitrary in the sense we want only positions outside the core. Also, note that the defect is placed on the CuO bonds here opposed to the center of the plaquette as was done in the main text for the LDOS. The spectral functions corresponding to these two situations are equivalent, meaning the energetics are very similar, but the shape of the $dI/dV$ curves will certainly change character since the gap amplitude takes on different values on the lattice sites.
@@
Both sets of curves, to the right and to the left of the vortex core, have a periodicity of four lattice spacings, just like that of the pure PDW in the text (the same conclusion holds for the placement of the defect at the center of the plaquette). However, unlike a pure LO state the data appears to have a relative shift of two lattice spacings when comparing the curves from the right to the left (again this holds true for the half-vortex placed at the center of the plaquette). As was discussed in the text, this is due to the jump in $\theta_{+}$ by $\pi/2$ across the half-vortex’s core and can be explained as an accumulation in phase due to the non-zero COM-momentum of our Cooper pairs; that is, $\mathbf{Q}\cdot\mathbf{r}=\pi/2$ if $\mathbf{r}=2a_{0}\mathbf{e}_{x}$. This is a smoking gun signature of the half-vortex since a bona-fide LO state would not break inversion symmetry like so.
@@
Now we compare the tunneling DOS of our topological defects inside and outside the core of the vortex [Figs. 15a and 15b, respectively]. The Abrikosov vortex and the double dislocation both resemble a free particle tunneling DOS, albeit some additional wiggles. This is simply because we are not right at the center of the core. Also, notice that the curves for these two defects are not in perfect agreement here, due to the form of the topological defect. Indeed, the electron DOS has to be distinct in the core of each vortex due to the difference in the respective defect. The half-vortex still possesses signatures of superconductivity in the core due to the fact it has FF character there. In fact, it resembles a squeezed in FF state.
@@
Outside the core the double dislocation and the Abrikosov vortex look almost identical to one another and also the pure PDW (LO state). This is because the topological defect has less of an influence on the electronic states far away from the core. The half-vortex has additional wiggles in its tunneling DOS corresponding to satellite peaks. These additional peaks arise here and not for the other two defects because of the half-vortex’s inequivalent Fourier components.
@@
Figure 16: Comparison plots of various Fermi surfaces. First, we provide a plot for (a) a defect-free PDW with an $s$-wave form factor compared with (b) a defect-free PDW with a $d$-wave form factor superimposed with the normal state dispersion (shown in light gray). The main arcs seen in both plots correspond to the normal state dispersion. In (c) we have a log plot of the Abrikosov vortex with a $d$-wave form factor, which greatly resembles the double dislocation and the pure PDW. Both defects prompt a slight redistribution of spectral weight, but nothing too dramatic like the half-vortex. Weaker features are uncovered with the log plots where the additional arcs correspond to the normal state dispersion shifted by ±$\pm\mathbf{Q}$.
@@
In this appendix we compare some additional spectral functions not included in the main text. First, let’s focus on the spectral functions corresponding to defect-free PDW states, one with an $s$-wave form factor and the other $d$-wave [Figs. 16a and 16b, respectively]. The normal state dispersion is retained in both cases along the so-called Fermi arcs, so certain features are robust. The LDOS looks more or less the same for both form factors (not shown). Figures 16c is the natural log of the spectral function corresponding to the Abrikosov vortex, which closely resembles the double dislocation (not shown). Slight redistributions of the spectral weights, occur for both these defects, but nothing as dramatic as that seen in the case of the half-vortex.